OpenStudy (anonymous):

Find 3 consecutive even numbers where the product of the smaller two numbers is 64 less than the square of the largest number. @johnweldon1993 @Jhannybean

OpenStudy (johnweldon1993):

hmm...well we only want even numbers...so lets let the first number be 2x...and we want consecutive even numbers...so we can do 2x + 2 and 2x + 4 Now...the product of the smaller 2 is 64 less than the square of the largest number \[2x \times (2x + 2) = (2x + 4)^2 - 64\] \[4x^2 + 4x = 4x^2 + 16x - 48\] \[-12x = -48\] \[x = 4\] so if 'x' is equal to 4...we can go back to the original 2x , 2x + 2, 2x + 4 thing...we know x = 4 so 2(4), 2(4) + 2, 2(4) + 4 our numbers are 8, 10, 12

OpenStudy (johnweldon1993):

now we can check....product of 2 smaller 8 X 10 = 80 is 64 less than the square of the bigger 12² = 144 144 - 64 = 80 so they work